OpenStudy (anonymous):

If y varies directly as x and y=6 when x=18, then find x when y=10.

OpenStudy (anonymous):

some of you tutors make me so angry...

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

Do you know how he got that answer? Do you know if he's right or not? How can you check?

OpenStudy (anonymous):

Hrm. Maybe the answer is 42

OpenStudy (anonymous):

Nope. It's a multiple choice question ;)

OpenStudy (anonymous):

Maybe it's one of the other answers.

OpenStudy (anonymous):

\[y = kx \]

OpenStudy (anonymous):

k= y/x =>6/18

OpenStudy (anonymous):

again y=kx x=y/k

OpenStudy (anonymous):

x= 10 * 18/6

OpenStudy (anonymous):

x= 30.... Now u know my method :)

OpenStudy (anonymous):

They've stopped paying attention Ishaan. You gave them the answer already. Why should they care how to do it again? If they have another question like this, they'll just ask you how to do it again.

OpenStudy (anonymous):

but how did u get 42 ... polpak

OpenStudy (anonymous):

but all they wnt are answers most of them

OpenStudy (anonymous):

Thank you for showing your work ishaan94. Greatly appreciated :)

OpenStudy (anonymous):

You can't always get what you want... You can't always get what you want... But if you try sometimes.. You just might find.. You get what you need!

OpenStudy (anonymous):

Of course they just want answers. That doesn't mean you should just give answers.

OpenStudy (anonymous):

ok i promise i'l try to explain frm next time :)

OpenStudy (anonymous):

I'm not sure if you noticed but he showed his work after you started complaining.

OpenStudy (anonymous):

its okay mathsux4real ur a good guy

OpenStudy (anonymous):

polpak is too

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Girl you mean (haha)

OpenStudy (anonymous):

but still polpak how did u get 42...

OpenStudy (anonymous):

yea sry good grl

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Yes, my frustration is in the fact that it only happened afterward. AND with the fact that if I asked you now that y varies directly with x and when y=15 x = 3 what will y be when x is 5. I'm not sure you'd understand how to do it..

OpenStudy (anonymous):

Tutoring is not about giving answers. It's about asking questions.

OpenStudy (anonymous):

Stop that.

OpenStudy (anonymous):

sry i thought u were asking me ... sry again

OpenStudy (anonymous):

Who are you talking to now polpak? Stop what?

OpenStudy (anonymous):

I was talking to Ishaan.

OpenStudy (anonymous):

I've been talking to Ishaan this whole time.

OpenStudy (anonymous):

Except when I asked if you could solve that last problem.